# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ X-A4 Cognitive Radio Networks (CRNs)
[82] Flexible integration of DNN with channel coding for iteration gain Turbo-style Multi-Carrier AE OFDM systems Substantial performance gain over conventional OFDM Complexity in soft-decision integration with channel decoding
[104] Variational AE-enhanced DNN for MIMO signal detection VAE-DNN-Det Simulated wireless networks 1 dB SNR gain over traditional methods Performance depends on channel condition spectrum
[106] Joint channel estimation, feature extraction, and symbol detection with AEs AE-based DNN Simulated MIMO channels Significant BER gains in high channel correlation scenarios Complexity of training deep networks for MIMO systems
[108] Embedded singular vectors in DAE for enhanced BER SVD-embedded DAE Simulated MIMO systems BER improvement by 10-18x over linear precoding at SNR of 10dB Limited to linear precoding scenarios
[102] Adaptive AE for dynamic interference management in beyond-5G systems ADL AE Multi-user Gaussian interference channels 2 dB improvement in BER over conventional schemes Performance highly dependent on interference estimation
[111] Efficient CSI feedback with robustness to noise PRVNet (VAE-based) Indoor and outdoor FDD-MIMO scenarios NMSE: $-27.7$ dB (indoor), $-13.9$ dB (outdoor) at CR=1/4 Limited to feedback channels with specific compression ratios
[113] Joint quantization and compression for CSI feedback AE with entropy bottleneck layer Indoor (5.3 GHz) and outdoor (300 MHz) CSI feedback NMSE: $-24$ dB (indoor), $-13$ dB (outdoor) at BPP=3.0 Requires specific tuning for optimal entropy bottleneck performance
[115] Joint optimisation of RIS and transmitter-receiver pairs RIS-aided AE (Encoder-Decoder with RIS) Simulated scenarios with obstruction losses SER: 18 dB gain at $l_{o}=10$ dB, SER=$10^{-5}$ Limited evaluation in scenarios with dynamic environmental changes
[116] Joint transceiver and ADC design optimization AE with optimized ADCs MIMO Rayleigh fading channels 12 dB gain at BLER=$10^{-2}$, 1 dB loss at BLER $\leq 10^{-3}$ Requires accurate ADC modeling for optimal results
[117] Multi-objective optimization for PAPR, BER, and spectral performance Convolutional AE Simulated 16-QAM, 4×4 MIMO configurations PAPR reduction: 4 dB at CCDF=$10^{-3}$ Computational overhead for multi-objective optimization
[118] Energy-based AE for multiuser noncoherent systems NC-EA Multicarrier SIMO fading channels BLER=$10^{-3}$ with 8 dB gain (single-user), 6.5 dB (multi-user) Performance depends on accurate sub-carrier power allocation
[119] Improved target detection accuracy over conventional CFAR AE replacing CFAR FMCW radar system simulations Correlation coefficient $\geq 0.73$ with actual target data Limited to automotive radar systems
[121] Improved BER and SNR performance for MIMO detection DNN-Det-VAE MIMO setups (64-QAM, 256-QAM) SNR gain: 1 dB at BER=$10^{-5}$ (4×4 MIMO, 64-QAM) Limited gains in high SNR conditions compared to advanced methods
X-A5 MIMO
In the realm of modern wireless communication, MIMO technology has emerged as the cornerstone for achieving higher data rates, improved spectral efficiency and reliability. As demand for high-performance systems grows with the advent of 5G and beyond, traditional signal processing techniques often fail to address the complexity and dynamic nature of MIMO systems. This has paved the way for the integration of advanced deep learning methodologies, particularly deep AEs. Deep AEs, known for their ability to extract meaningful representations and compress high-dimensional data, offer a promising solution to tackle various challenges in MIMO systems. By treating the communication process as an end-to-end optimisation task, AEs can jointly optimise transceiver design, CSI feedback, and interference mitigation. This approach not only simplifies system architectures but also significantly enhances key performance metrics such as BER, spectral efficiency, and energy consumption. The application of deep AEs spans a wide range of MIMO functionalities, including hybrid beamforming, reduction of PAPR, and robust detection under adverse channel conditions.
Signal Detection and Processing
Several recent studies have explored the use of AE-based frameworks and deep learning techniques to improve signal detection and processing in MIMO systems. Bui et al. [103] introduced two innovative systems, BWAE-MIMO and SWAE-MIMO, which use AE to improve signal detection performance. Among these, the BWAE-MIMO system, equipped with the radio transformation network (RTN) detector, achieved BER performance comparable to the maximum likelihood detection (MLD) approach in 2×2 MIMO configurations. This highlights the potential of AEs to compete with traditional methods while leveraging the advantages of deep learning. Building on this concept, Omondi and Olwal [121] proposed the DNN-Det-VAE architecture, which incorporated VAEs into a deep learning framework for MIMO detection. Their method demonstrated a SNR gain of 1 dB in 4×4 MIMO setups with 64-QAM and 256-QAM constellation formates. The approach exhibited superior BER performance compared to traditional methods such as zero force (ZF) and minimum MSE, which demonstrated its effectiveness under challenging channel conditions. In a similar vein, Zhao et al. [107] developed a VAE-based signal detection method adapted to MIMO-OFDM-IM systems. Their approach achieved BER performance that was not only close to existing hand-crafted detection algorithms but also featured reduced runtime, thus underscored the efficiency of deep learning-driven solutions. Kavitha et al. [106] explored the application of AE-based DNN for MIMO signal processing. Using the dimensionality reduction capabilities of AEs, their framework demonstrated robustness to channel impairments and delivered significant performance improvements in BER and spectral efficiency compared to conventional MIMO detection techniques. This study also highlighted the scalability and flexibility of deep learning approaches to address complex MIMO communication challenges.
Hybrid Beamforming and Advanced Architectures
Tao et al. [109] introduced an innovative AE-based hybrid beamforming method (AE-HB) for mmWave. By mapping the hybrid beamforming problem to an AE neural network, their approach transformed the traditionally non-convex optimisation problem into a neural network training process. This method demonstrated a significant 3 dB improvement in BER compared to conventional linear beamforming systems, thus showcased the potential of using neural networks to optimise hybrid beamforming in next-generation communication systems. Building on the integration of deep learning into MIMO systems, Cherif et al. [110] developed an end-to-end learning approach aimed at mitigating nonlinear distortions, such as those introduced by high-power amplifiers in massive MIMO systems. Their framework used deep neural networks and AEs to jointly optimise the transmitter and receiver under Rayleigh fading conditions. The proposed approach achieved near-linear SER performance at high SNRs with 16-QAM modulation, which significantly outperformed traditional distortion mitigation techniques. Together, these studies highlight the transformative potential of AE-based methods to address complex challenges in MIMO communication, including beamforming and nonlinearity compensation, while improving system reliability and efficiency.
Channel State Information (CSI) Feedback
Jang et al. [105], Hussien et al. [111], and Ravula et al. [113] each addressed the challenge of CSI-efficient feedback in MIMO systems through innovative deep learning approaches. The first proposed a deep AE-based method that significantly reduced normalised MSE and demonstrated robust performance under feedback errors and delays, which ensured reliable CSI reconstruction. Building on this, the later introduced PRVNet, a VAE-based model optimised for high compression scenarios. PRVNet achieved exceptional NMSE values of $-27.7$ dB in indoor environments and $-13.9$ dB in outdoor settings in a compression ratio of 1/4, which outperformed many existing methods. Similarly, the latter proposed an entropy bottleneck layer within an AE-based feedback framework, which enabled joint optimisation of bit rate and distortion. This approach achieved NMSE values of $-24$ dB for indoor channels and $-13$ dB for outdoor channels at specific bit-per-pixel (BPP) values, which demonstrated superior reconstruction quality over traditional methods. Collectively, these studies highlighted the potential of AE and VAE-based techniques to improve CSI feedback efficiency and accuracy in diverse wireless environments.
Interference Mitigation and Robustness
Wu et al. [102] and Sahay et al. [114] explored innovative deep learning approaches to enhance interference mitigation and system robustness in wireless communication. The first proposed an adaptive AE designed for 5G and beyond systems, which effectively addressed dynamic interference challenges. Their method demonstrated robust BER performance, particularly in strong and very strong interference scenarios, which demonstrated adaptability to varying network conditions. The second, on the other hand, introduced a denoising AE aimed at countering adversarial attacks in massive MIMO power allocation systems. This approach was shown to be effective in various attack models, ensuring resilience and maintaining system performance even under adversarial conditions. Together, these studies highlighted the critical role of AE-based techniques in mitigating interference and enhancing the robustness of next-generation wireless systems.
Performance Optimisation for End-to-End Communication
Zhang et al. [108] and Van et al. [118] focused on optimising end-to-end communication performance in wireless systems using advanced deep learning frameworks. Zhang et al. introduced a singular value decomposition (SVD)-embedded deep AE for MIMO systems, which demonstrated significant BER improvements. Their approach achieved up to an 18-fold reduction in BER compared to traditional precoding methods at an SNR of 10 dB, which highlighted the potential for embedding mathematical transformations within AE structures. Van et al. developed a noncoherent energy AE (NC-EA) system for single-input multiple-output (SIMO) setups, which addressed both single-user and multi-user scenarios. The proposed NC-EA system achieved substantial SNR gains, including 8 dB for single-user configurations and 6.5 dB for multiuser systems at a BLER of $10^{-3}$. These advancements underscore the effectiveness of integrating deep learning into end-to-end communication design to achieve superior reliability and efficiency.
Specialized Applications and Novel Methods
Shin et al. [112] introduced the amplitude phase quantisation with neural networks (APQNN) framework for relay systems, which achieved significant BER improvements. Their approach demonstrated performance close to $10^{-3}$ BER at high SNR levels, which emphasised its effectiveness in the optimisation of the relay system. Similarly, Erpek et al. [115] developed a RIS-aided AE system for joint transmitter-receiver optimisation. This method achieved substantial reductions in SER, with performance gains of up to 18 dB under high obstruction losses, which highlighted the potential of RIS in improving end-to-end communication efficiency. Ye et al. [116] proposed an AE-based MIMO communication framework that incorporates end-to-end optimised analogue-to-digital converters (ADC). This approach resulted in a gain of 12 dB at a BLER of $10^{-2}$, closely matching the unquantified lower bound baselines. In another innovative effort, Huleihel et al. (2024) [117] introduced a convolutional AE framework for PAPR reduction in MIMO-OFDM systems. Their method achieved a 4 dB reduction in a complementary cumulative distribution function (CCDF) of $10^{-3}$, demonstrating effectiveness in optimising spectral and detection performance simultaneously. Further extending the applications of AEs, Kang et al. [119] replaced traditional constant false alarm rate (CFAR) algorithms with an AE-based approach in frequency-modulated continuous wave (FMCW) radar systems. The proposed method achieved a correlation coefficient of 0.73 or higher in the target detection results, which highlighted its robustness in signal processing tasks. Furthermore, Zhang et al. [120] proposed a deep AE-based architecture for Z interference channels. This innovative system achieved up to two orders of magnitude reduction in BER, particularly under imperfect CSI conditions of the channel, which demonstrated its resilience to interference and CSI estimation errors.
@@ X-A5 MIMO
TABLE XI: Comparison of AE-based models in Optical Communication
@@
[122] Improved BER Dense (ReLU) Binary vectors, AWGN Better than half-rate 16-QAM Idealized communication assumptions
@@
[124] Train in real channels, power control VAE, Feedforward AWGN, AIL, AICN, RBF Faster training, lower BLER Challenges at high dimensions
@@
[129] Adaptation to nonlinear fibre links Dense (ELU) AWGN, CD, KNL High spectral efficiency gain Training complexity
[130] Amplification noise vs. nonlinearity Dense (ReLU) Differentiable fibre channel 0.13 bit/4D improvement Model discrepancy
[137] Optimise constellation shaping Bitwise Synthetic datasets, Wiener phase noise Improves 64-ary performance Lack of detailed complexity analysis
[138] Differentiable CPE Fully-connected (ReLU) Simulated signals, Wiener phase noise µ = 4 provides best performance Additional computational complexity
[139] Differentiable channel model Dense (ReLU) AWGN, RBF, simplified fibre-optic State-of-the-art hardware Relies on accurate channel models
[140] Design constellations for nonlinearities Dense (SELU) Nonlinear fibre channel, ASE noise 0.13 bit/4D gain Model agnostic to embedded channel
[141] Optimise transceiver chain Fully-connected (ReLU) Labelled input-output pairs 42 Gb/s at $<$ 6.7% HD-FEC Sensitivity to distance variations
[142] Optimise input probability distributions Distance-agnostic AE Experimental data, Generative model Outperforms uniform distributions Optimisation on specific model assumptions
[145] Optimise geometric constellations Dense (ReLU) fibre channel with WDM Up to 0.2 bits/symbol GMI gain Implementation penalty
X-B Optical Communications
The existing research focussed on the application of deep learning techniques, particularly the use of AEs and VAEs in various optical communication systems, addressing unique challenges and proposing innovative solutions.
In the field of OWC, Si-Ma et al. [123] explored the use of AEs for the design of the SISO transceiver operating in a Poisson channel. The proposed AE is specially designed for the poison channel with the help of the square root (SR) transform. This shows a unique adaptation rather than a standard architecture. Raj et al. [124] claimed that classical AEs are not inherently designed to handle noisy / disrupted received signals, and some methods used in the normalisation layer to enforce power constraints can restrict the ability of the models to find the optimal solution. Therefore, they emphasised the adaptation of end-to-end learning frameworks to handle noise corruption in transmitting symbols using VAE. VAEs are better suited for handling noisy latent representation. In addition, integrating domain knowledge into the training process accelerates model performance and efficiency. A modified los function of CE has been utilised to minimise the transmit power while maximising the reconstruction likehood. The study extended beyond the AWGN model to include various noise scenarios, such as additive independent Laplace noise (AILN) and Additive independent cauchy noise (AICN), along with RBF.
However, Kim et al. [125] developed a versatile AE framework that can be adapted to arbitrary coding configurations, thus improving the scalability of the transceiver design and the performance of the VLC. The work addressed fixed AE computational structures that lack flexibility in varying the lengths of message bits and codewords. In the same area of VLC, Mohamed et al. [26] aimed to address the problem of high PAPR by proposing a long-short-term memory-AE (LSTM-AE) model. This model used for the first time the LSTM-AE architecture with a fully connected layer using the LeakyReLU activation function. It shows the ability to maintain BER while effectively reducing PAPR, showcasing its superiority over existing techniques. Furthermore, Zhai et al. [126] research delved into UOC, showcasing the superior performance of AE models in handling channel attenuation due to water scattering and absorption.
The enhancement of FSO communication systems using AEs has recently been investigated by Zahng et al. [127] and Cao et al. [128]. The first used an end-to-end AE-based transceiver that uses DNN to enhance UAV-to-ground FSO communication, particularly under strong atmospheric turbulence. The proposed model demonstrated a significant improvement in BER performance and transmission efficiency under the Rytov variance channel compared to traditional methods such as pulse position modulation (PPM). The later investigated convolutional AEs with ELU activation function under unknown CSI scenarios. The proposed model outperformed conventional maximum likelihood (ML) methods when CSI is not available in the atmospheric turbulence channel (Gamma-Gamma distribution).

### v2

@@ VI-B1 Dispersion and Nonlinearity
[114] optimize constellation shaping Bitwise Synthetic datasets, Wiener phase noise Improves 64-ary performance Lack of detailed complexity analysis
[113] optimize input probability distributions Distance-agnostic AE Experimental data, Generative model Outperforms uniform distributions Optimisation on specific model assumptions
[108] optimize transceiver chain Fully-connected (ReLU) Labelled input-output pairs 42 Gb/s at << 6.7% HD-FEC Sensitivity to distance variations
[112] Adaptation to nonlinear fiber links Dense (ELU) AWGN, CD, KNL High spectral efficiency gain Training complexity
[117] Design constellations for nonlinearities Dense (SELU) Nonlinear fiber channel, ASE noise 0.13 bit/4D gain Model agnostic to embedded channel
[118] optimize geometric constellations Dense (ReLU) fiber channel with WDM Up to 0.2 bits/symbol GMI gain Implementation penalty
[120] Differentiable CPE Fully-connected (ReLU) Simulated signals, Wiener phase noise $\mu$ = 4 provides best performance Additional computational complexity
VI-B2 Equalization and Carrier Recovery
Achieving robust equalization and carrier recovery in coherent optical fiber communication systems remains a critical and complex challenge. Traditional algorithms such as the adaptive zero-forcing (ZF) approach suffer from convergence issues during startup, while independent signal analysis (ISA) methods are prone to noise-induced numerical instabilities [121, 24]. On the other hand, bootstrapping blind equalization algorithms, especially in systems using adaptive equalizers, has proven difficult, with commonly employed techniques such as the constant modulus algorithm (CMA) and its variants failing to converge at critical operating points [111]. These issues are further exacerbated in systems that use higher-order modulation formats and PCS, where equalizers face significant difficulty maintaining reliable performance [111, 115]. Conventional decision-directed algorithms also struggle in low SNR regimes, while the convergence of CMA becomes notably slower for complex modulation schemes [116]. Furthermore, designing blind receivers for PCS-modulated systems poses additional complexity, particularly in time-varying channels where ISI and noise further degrade performance [115, 114]. Despite extensive research, there remains a lack of effective blind maximum likelihood (ML)-based equalizers suitable for PCS formats and dynamic conditions [115].
In response to these challenges, pilot-less VAE equalizers demonstrated substantial gains in channel estimation accuracy, which can eliminate the need for overhead while maintaining high performance [121]. Frequency-domain blind equalization schemes and vector quantized VAE (VQ-VAE) equalizers have also been proposed to deliver lower complexity without sacrificing accuracy [116]. However, VQ-VAE equalizers require additional computation to reconstruct channel observations from decoded symbols. In addition, training VQ-VAE equalizers may face difficulties in computing the ELBO for certain channels.
On the other hand, VAE-based linear equalizers (VAE-LE) have been explored for blind equalization and channel estimation. Lauinger et al. [115] investigated adaptive blind equalizers based on variational inference for carrier recovery in optical communications, specifically leveraging VAEs for channel estimation. A dispersive linear optical dual-polarization channel is employed, along with an AWGN and ISI time-varying channel. They demonstrated that VAEs can outperform traditional methods such as CMA in fixed and time-varying channels. In the same direction, they addressed the improvement in the performance of blind adaptive equalizer bootstrapping, specifically addressing the challenges faced by algorithms such as CMA during the start-up at critical working points [111]. Their work demonstrated that novel VAE-based equalizers could achieve reliable convergence at these critical points for both standard and PCS formats, unlike CMA and its variants which often failed. The demonstration used a fiber channel using a linear frequency domain channel matrix and assuming a time-invariant channel during the bootstrapping phase AWGN and first-order PMD. Critically, this work mainly focused on linear impairments and assuming that potential nonlinearities were negligible or compensated separately. The extension of VAE-LE towards nonlinearities remains an area for future research. Furthermore, the performance of the VAE-LE can be significantly influenced by hyperparameters such as the symbol rate and the filter length, which might lead to performance penalties under certain conditions.
Alternatively, Song et al. [116] presented a novel frequency-domain blind equalization scheme using VQ-VAEs. It demonstrated that this approach can achieve performance similar to that of time-domain methods while reducing computational complexity, particularly in long-memory polarization division multiplexing (PDM) channels. This implies that the VQ-VAE equalizer can overcome the slow convergence of traditional blind equalization methods, such as CMA, especially for high-order modulation formats. However, there are challenges in the training process, as the ELBO computation for VAE equalizer training may face difficulties for certain channels. Additionally, the VQ-VAE equalizer requires additional computation to reconstruct channel observations from decoded symbols.
VI-B3 Constellation Shaping
Constellation shaping plays an essential role in modern optical fiber communication systems. It offers the potential to enhance mutual information (MI), spectral efficiency, and transmission reach. However, both geometric shaping (GS) and probabilistic shaping (PS) present non-trivial design and implementation challenges. In GS, non-Gray labeling incurs implementation penalties and complicates receiver design, while PS typically requires a dedicated shaping encoder, which poses difficulties for efficient and parallelized implementation [118]. The overall complexity of shaping, including its iterative demapping requirements and reliance on non-binary FEC, adds further constraints to practical deployment. A key challenge in both shaping paradigms lies in jointly optimizing the constellation shape and system parameters to adapt to real-world impairments. These include nonlinear channel effects, fiber dispersion, and hardware imperfections, all of which must be considered during optimization to achieve shaping gains in practical fiber-optic channels [117].
To address these multifaceted challenges, AE-based solutions have emerged as promising solutions. These models enable joint geometric and probabilistic shaping through AE end-to-end learning frameworks that adaptively optimize constellation layouts and symbol distributions for nonlinear and dispersive fiber channels [117, 118, 58]. In probabilistic shaping with AE also poses some challenges, such as it requires a shaping encoder, which can complicate parallelized implementation, Sampling uniformly from constellation points complicates the process for channels with memory, such as optical fiber, and difficulties in estimating gradients of cost functions related to constellation probabilities. Innovative training techniques now enable AEs to sample from constellation distributions without approximation to improve both shaping accuracy and training stability [57].
Furthermore, a series of recent works have proposed differentiated bit-level probabilistic shaping (BPS) modules that enable flexible and optimized constellation generation within deep learning pipelines [119, 114]. These approaches outperform conventional shaping methods by achieving significant gains in MI and transmission reach, particularly under nonlinear channel conditions. However, there are open challenges in geometric shaping with AE such as it incurs an implementation penalty due to non-Gray labeling in conventional BICM systems and it requires iterative demapping or non-binary FEC. The works in [120] investigated the optimization of both geometric and probabilistic constellation shaping within Wiener phase noise channels. The core objective is to leverage deep learning AE to design constellations and associated transceiver functions that are robust to phase noise by integrating CPE directly into the end-to-end optimization process. This was particularly explored using the BPS algorithm, and also variations of the Viterbi & Viterbi (V&V) algorithm. However, this integration required the development of a differentiated implementation of BPS/CPE. The standard BPS algorithm is inherently non-differentiable due to its reliance on operations like arg min, which prevents the direct application of gradient descent necessary for end-to-end deep learning optimization.
A recent work in [122] introduced a new approach for PCS in communication systems by leveraging automatic differentiation and importance sampling. The importance sampling in this work is used to draw symbols from a simplified distribution while continuously updating the approximation to the actual constellation probabilities. This method leads to a more stable and straightforward optimization process, as the gradients of the cost functions with respect to the constellation point probabilities become exact, removing potential errors caused by manual gradient correction. However, the proposed method is sensitive to how it is started and requires careful setting of various parameters. Additionally, it demands precise adjustments of hyperparameters to avoid getting trapped in local minima, which means that poor choices in setup could lead to suboptimal performance. Furthermore, relying on standard automatic differentiation without additional correction steps makes the system simpler, but it also carries the risk that the optimization process may not explore the entire solution space effectively.
While joint geometric and probabilistic shaping through AE-based systems provides a transformative direction for the design of next-generation fiber optic networks capable of supporting high-throughput, low-latency, and energy-efficient communications, the complexity of jointly learning probabilistic and geometric shaping poses a significant hurdle in optimizing the physical layer. Therefore, further exploration is needed in the joint learning of probabilistic and geometric shaping, particularly in determining the optimal placement of constellation points. Generally, learned constellations are often tailored for specific receiver algorithms, making it challenging to achieve optimal performance universally.
VI-C Optical Wireless Communications (OWC)
AEs are being actively explored to address specific challenges within various optical wireless communication systems, including free-space optical (FSO), visible light communication (VLC), underwater optical communication, where those systems face challenges related to the unique properties of light as a transmission medium, including atmospheric effects and intensity/nonnegativity constraints, as well as general challenges in applying deep learning techniques. as shown in Fig. 8.
Optical Wireless CommunicationSolutionsChallenges Channel Characteristics and Modeling: non-negativity, intensity modulation constraints, variable channel conditions FSO Specific Challenges: coding efficiency, degradation in strong turbulence Underwater Optical Communication: severe channel attenuation, hardware design and signal processing AE-based OWC systems, sparse-representation AE, … active challenge area AE-based transceiver for FSO, convolutional AEs … active challenge area AE-based model tailored to underwater optical communication … active challenge area
Figure 8: Optical wireless communication challenges and solutions
@@ VI-C Optical Wireless Communications (OWC)
Table XI lists the state-of-the-art comparison of AE-based models for optical wireless communication.
TABLE XI: Comparison of AE-based models in optical wireless communication
@@
[21] optimize constellation for near Shannon capacity Dense (ReLU), softmax Shared communication, AWGN Improves SER over TDMA/FDMA computational complexity
@@
[122] new cost functions and a weight learning method NN-layers, COCOB Optimizer AWGN, 256-QAM MI differences at maximum around 0.01 bits per symbol Sensitive to hyperparameters, Local Minima Risk
@@
VI-C1 Channel Characteristics and Modeling
OWC systems face distinct challenges arising from the physical nature of the optical channel, including nonnegativity and intensity modulation constraints, as well as variable channel conditions that differ markedly from traditional radio frequency (RF) models [124, 125]. Furthermore, developing effective communication techniques that do not rely heavily on existing analytical channel models remains a pressing challenge, as these models often fail to capture the complexity of real-world OWC channels [125]. These difficulties are particularly pronounced in Poisson channels, which are commonly used to model photon-limited optical intensity channels.
To address these challenges, AE-based OWC systems have been shown to outperform conventional QAM formats in terms of BER, particularly under multipath fading conditions [124]. They also present a viable solution to the modeling and design challenges posed by Poisson channels. AE end-to-end learning techniques tailored specifically for these channels enable transceiver architectures to learn the unique statistical characteristics of photon-limited communication without requiring precise analytical models [21]. Recent proposals demonstrate that sparse-representation AE (SR-AE) can match or exceed the performance of traditional transceivers while offering reduced computational complexity [126]. They addressed the challenge of designing a transceiver over the Poisson channel, which often relies on approximations and numerical searches. The AE architecture may not fully meet the requirements for non-negative and intensity constraints.
Another open issue is the unclear impact of varying the modulation order on overall system performance. Both high and low-order modulation schemes can hinder communication efficiency under different OWC conditions. This requires adaptive modulation strategies that can dynamically respond to channel variability [124]. Traditional techniques such as Quadrature Amplitude Modulation (QAM) are often not suitable for OWC environments, especially in the presence of multipath fading, which leads to signal distortion and degraded performance [124].
VI-C2 VLC Systems
In the area of VLC, Mohamed et al. [36] aimed to address the problem of high PAPR by proposing a long-short-term memory-AE (LSTM-AE) model. This model used for the first time the LSTM-AE architecture with a fully connected layer using the LeakyReLU activation function. It shows the ability to maintain BER while effectively reducing PAPR with its superiority over existing techniques. However, increasing the hyperparameter beyond 0.3 does not significantly improve the PAPR values. Additionally, while the LSTM-AE has fewer parameters compared to a Dense AE, its processing time was slightly longer than that of the Dense-AE, which is a consideration in real-time applications.
On the other hand, Kim et al. [123] developed a versatile AE framework known as the autoencoding graph neural network (AEGNN) to address limitations in existing AE models and improve scalability of the transceiver design, particularly in VLC. The work addressed fixed AE computational structures that lack flexibility in varying the lengths of message bits and codewords. The study demonstrated that AEGNN can maintain an almost identical SER to traditional methods for different coding configurations. While the framework demonstrates versatility for various configurations, the limitation in applying AE architectures to higher dimensions remains. The curse of dimensionality is even an insurmountable barrier for AE architecture in high-dimensional scenarios. This suggests that while AEGNN offers flexibility for different configuration pairs within certain limits, scalability to much greater than 8 bits remains a significant challenge for current AE architectures.
For probabilistic shaping in VLC, training a sampling mechanism for symbols drawn from a finite set using ML is challenging [129, 57]. There are difficulties in estimating the gradients of cost functions related to the probabilities of constellations.
VI-C3 FSO Systems
An end-to-end AE-based transceiver has been proposed specifically for FSO communication systems [127]. The research aims to address the challenges posed by atmospheric conditions in UAV-to-ground FSO links, where atmospheric turbulence can cause intensity fluctuations and beam wandering. Notably, the proposed model demonstrated a remarkable 12 dB performance gain over traditional pulse position modulation (PPM) transmitters, especially in the presence of atmospheric distortions such as scintillation and beam spread. However, maintaining coding efficiency while increasing transmission power directly impacts the overall effectiveness of transmission, which is an essential requirement for reliable long-distance optical links. Additionally, there is a challenge in optimizing transmission schemes due to a trade-off between symbol differentiation and error floor. In addition, performance degradation in strong turbulence is still an open challenge.
On the other hand, Cao et al. [128] investigated the use of convolutional AEs for improving FSO communication through end-to-end AE learning. Their work focused on addressing complex problems of practical FSO deployments due to factors like sensitivity, pointing inaccuracies, and atmospheric turbulence. The proposed model demonstrated improved performance compared to conventional ML detection or estimation methods when CSI is unknown. While the proposed system shows improved resilience against turbulence compared to conventional systems, its BER performance still degrades with increasing turbulence intensity. This highlights that despite the improvements, the system remains vulnerable to strong atmospheric effects, which emphasizes the need for effective modeling and mitigation strategies.
VI-C4 Underwater Optical Communication Systems
Underwater optical communication presents a unique set of challenges, most notably severe channel attenuation caused by scattering and absorption effects in the aquatic environment [18]. These impairments significantly degrade signal strength over distance. In addition, conventional transceiver technologies often fail to mitigate these losses. Consequently, there is a pressing need for innovation in both hardware design and signal processing approaches. Achieving high-bandwidth communication rates in such conditions remains particularly difficult, further constraining the deployment of real-time and data-intensive underwater applications.
In response to these limitations, recent work has proposed an AE-based model tailored to underwater optical communication. This data-driven architecture has demonstrated the potential to enhance transceiver performance, particularly under severe channel attenuation conditions [18]. Importantly, the AE structure exhibits optimal performance in scenarios involving high symbol complexity, where traditional modulation and detection schemes often struggle. By jointly learning an end-to-end mapping between transmitted and received signals, the AE effectively adapts to the statistical characteristics of the underwater channel.
However, the proposed AE architecture does not fully meet the requirements for non-negative and peak-limited optical communication. These physical constraints necessitate modifications to the standard AE output to be compatible with optical transmission hardware, which is a crucial consideration for practical implementation.
VI-D Semantic Communication
Semantic communication focuses on transmitting the meaning of information, which introduces challenges beyond traditional reliable bit transmission, such as quantifying meaning, dealing with background knowledge differences, and handling semantic noise, as listed in Fig. 9.
Semantic CommunicationSolutionsChallenges Defining and Quantifying Meaning: define, model, and quantify meaning, high recovery rates Semantic Noise and Quality Metrics: lack of objective metrics to model and evaluate semantic noise Background Knowledge Alignment: common BK between communication nodes, convey AE-based systems to intended meaning AE-based technique for learning conceptual spaces , … active challenge area AE-based image quality metric … active challenge area AE-based semantic communication framework, … active challenge area
Figure 9: Semantic communication challenges and solutions
VI-D1 Defining and Quantifying Meaning
The question of how to formally define, model, and quantify meaning remains one of the most persistent challenges in artificial intelligence and cognitive science. Despite current research, there is no consensus on a unified framework for modeling meaning in a way that is both computationally tractable and semantically expressive. Existing approaches often struggle with interpretability and fail to provide a clear understanding of the learned representations, which can limit their utility in applications that require human-aligned semantic reasoning. The absence of a standardized model for capturing conceptual meaning impedes progress in areas such as explainable AI, semantic search, and cross-modal reasoning.
To address these gaps, recent work has proposed an AE-based technique for learning conceptual spaces which is a framework initially inspired by cognitive theories of concept representation [130]. This approach enables the unsupervised learning of interpretable, semantically structured domains from raw input data, guided by high-level property labels. While the proposed framework learns a domain from raw data, the challenge of obtaining semantic property labels for training remains difficult. The study primarily considered linear dimensions and Euclidean semantic distortion, thereby a need for exploration into more general approaches is required. Furthermore, the lack of interpretability in deep learning systems was a motivation, but the interpretability of the learned dimensions in complex scenarios beyond the tested datasets could still be challenging. Crucially, achieving high recovery rates in scenarios with minimal changes in data, such as time series data, is a significant problem [131]. This suggests that AEs may struggle to capture and reproduce subtle semantic nuances or small changes in meaning that do not drastically alter the underlying data structure.
VI-D2 Semantic Noise and Quality Metrics
In the context of image transmission, particularly in emerging 6G communication systems, the lack of objective metrics to model and evaluate semantic noise poses a significant challenge. Unlike traditional distortions such as compression artefacts or Gaussian noise, semantic noise pertains to degradations that alter the meaning or recognisability of visual content. Existing noise removal techniques are largely ineffective in addressing these deeper, meaning-altering perturbations, as they typically focus on pixel-level fidelity rather than semantic integrity [132]. Currently, there is no widely accepted quality metric capable of capturing and quantifying semantic noise. Conventional full-reference and no-reference quality metrics often do not account for the semantic content of images, instead relying on structural or perceptual similarity measures that do not reflect human understanding of image quality [132].
To address this gap, an AE-based image quality metric (AEQM) has been proposed to explicitly model and quantify semantic noise [132]. A convolutional AE architecture is utilized to handle the semantic content of the images. To quantify semantic noise, the model calculates the MSE between the latent vector spaces of undistorted and distorted images. The proposed model demonstrated a high correlation (88%) with subjective quality assessments. However, the AEQM requires considerable computational power due to its complex architecture, which is particularity less efficient than simpler metrics like PSNR. In addition, while showing high correlation, the unpredictability of no-reference quality metrics in general could still pose challenges.
VI-D3 Background Knowledge Alignment
In scenarios involving multiple nodes (like relay channels), there is a challenge in achieving common background knowledge (BK) between communication nodes [133]. Traditional relay schemes are often inadequate for semantic transmission, as they primarily focus on symbol-level forwarding and lack the mechanisms to account for contextual disparities between the source and destination [133]. This limitation becomes more pronounced in scenarios requiring the accurate transmission of semantic meaning, where the destination interpretation heavily relies on pre-existing knowledge that may not align with the source’s context [133].
To address these issues, an AE-based semantic communication framework has been proposed, specifically designed to operate over wireless relay channels [133]. This framework improves semantic understanding by jointly learning to encode and decode messages in a way that is robust to variations in background knowledge. A key component of this approach is the semantic forward (SF) protocol, which enables relay nodes to forward semantically meaningful representations rather than raw symbols, which can thereby bridge knowledge gaps between communicating parties. Although AE-based systems can encode and decode sentences from a semantic dimension, it is a challenge for AE-based systems to convey the intended meaning accurately when aligning the interpretation of this semantic information when BK varies between transmitter and receiver.
Table XII lists the state-of-the-art comparison of semantic communication and hybrid quantum-classical communication.